Hello again!
Another question :)
Hope you can answer and thank you very much if you did :)
The length of a triangle is twice its width. If its length is decreased by 3m and its width is increased by 2m, the area of the resulting triangle is 30m^2. Find the dimensions.

Answer by Cromlix(4381)

Hi there,
I think you mean 'the height of a
triangle is twice its base width'
If so:
Base width = x
Height = 2x
If height is decreased by 3m
2x - 3
If the base width is increased by 2m
x + 2
Area = 1/2( base width x height)
30 = 1/2( 2x - 3)(x + 2)
30 = 1/2(2x^2 + x - 6)
30 = x^2 + x/2 - 3
Form equation
(ax^2 + bx + c = 0)
x^2 + x/2 - 33 = 0
Multiply tro' by 2
2x^2 + x - 66 = 0
Factorise
(2x - 11)(x + 6) = 0
2x - 11 = 0
2x = 11
x = 5.5m
....
x + 6 = 0
x = -6 (Not real as -ve)
......
Therefore,
Length = 8m
Width = 7.5m
of new triangle.
Hope this helps:-)